# Model-Based Transfer Learning for Contextual Reinforcement Learning

### v3

@@ 4.1 Modeling assumptions
A linear function is used to model the generalization gap function, formally $\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x^{\prime})\simeq\theta|x-x^{\prime}|$, where $\theta$ is the slope of the linear function and $x$ and $x^{\prime}$ are the context of the source task and target task, respectively.
@@ 4.2 Bayesian optimization
Gaussian process (GP) regression. Within the framework of BO, we model the source training performance ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\forall x\in X\setminus x_{1:k}}$ using Gaussian process (GP) regression. Specifically, the function ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ is assumed to follow a GP $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)\sim\mathcal{GP}(\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)],k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})))$, where $k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ is the covariance function, representing the expected product of deviations of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{\tilde{x}},}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ from their respective means. Let $D_{k-1}$ denote the data observed up to iteration $k-1$, consisting of the pairs $\{(x_{i},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{i},x_{i})})\}_{i=1,...,k-1}$. The estimated performance $\hat{J}_{k}$ after querying $k-1$ samples is updated as more samples are obtained. The posterior prediction of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}_{k}}$ at a new point $x$, given the data $D_{k-1}$ and previous inputs ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}}$, is normally distributed as $P(\hat{J}_{k}\mid D_{k-1})=\mathcal{N}(\mu_{k}(x),\sigma_{k}^{2}(x))$. $\mu_{k}(x)$ and $\sigma_{k}^{2}(x)$ are defined as $\mu_{k}(x)=\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)]+\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{y}$ and $\sigma_{k}^{2}(x)=k(x,x)-\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{k}$, with $\mathbf{k}$ being the vector of covariances between $x$ and each $x_{i}$ in the observed data, i.e., $\mathbf{k}=[k(x,x_{1}),\ldots,k(x,x_{k-1})]$, and $\mathbf{K}$ is the covariance matrix for the observed inputs, defined as $\mathbf{K}=[k(x_{i},x_{j})]_{1\leq i,j\leq k-1}$. This enables the GP to update its beliefs about the posterior prediction with every new observation, progressively improving the estimation.
@@
$a(x;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}})=\mathbb{E}_{x^{\prime}\in X}[[\mu_{k-1}(x)+\beta_{k}^{1/2}\sigma_{k-1}(x)-\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})-{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{{1:k-1}},x^{\prime}})]_{+}]$ (4)
@@ 4.3 Regret analysis
We define the search space $X_{k}$ at iteration $k$ as a subset of $X$, with each element $x^{\prime}\in X_{k}$, such that $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1},x^{\prime}})\leq\hat{J}(\pi_{x_{k}},x^{\prime}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\Delta J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x_{k}},x^{\prime}})$.
@@
For a given $\delta^{\prime}\in(0,1)$ and scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$, the cumulative regret $R_{K}$ is bounded by $\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}$ with probability at least $1-\delta^{\prime}$.
@@ 5.1 Setup
Baselines. We consider two types of baselines when evaluating our proposed algorithm: canonical and multi-policy. The canonical baselines are selected to validate multi-policy training; the multi-policy training baselines are heuristic methods designed to validate the Bayesian optimization approach. The canonical baselines include: (1) Independent training, which involves independently training separate models on each task; and (2) Multi-task RL, where a single “universal” context-conditioned policy is trained for all tasks. The multi-policy baselines include: (3) Random selection, where each training task is chosen uniformly at randomly; (4) Equidistant strategy, which selects training tasks by equally subdividing the context space based on a given training budget, and then trains them in lexicographical order; (5) Greedy strategy, which greedily selects the next source task based on Assumption 1 and fixed training performance; and (6) Sequential Oracle transfer, which has access to generalized performance corresponding to policies for all tasks (including those not yet selected) and uses that information to greedily select the best source task at each step.
Proposed method. We evaluate MBTL with the scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$.
DRL algorithms and performance measure. We utilize Deep Q-Networks (DQN) for discrete action spaces [29] and Proximal Policy Optimization (PPO) for continuous action spaces [36]. For statistical reliability, we run each experiment three times with different random seeds. We evaluate our method by the average performance across all $N$ target tasks after training up to $K=$15 source tasks or the number of source tasks needed to achieve a certain level of performance. We employ min-max normalization of the rewards for each task, and we provide comprehensive details about our model in Appendix A.4.1.
@@ 5.2 Traffic benchmark experiments
Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
Traffic Signal Road Length 0.9409 0.8242 0.9366 0.9337 0.9349 0.9364 0.9409
Traffic Signal Inflow 0.8646 0.8319 0.8699 0.8712 0.8682 0.8432 0.8768
Traffic Signal Speed Limit 0.8857 0.6083 0.8872 0.8872 0.8874 0.8867 0.8876
Eco-Driving Penetration Rate 0.526 0.1945 0.6212 0.6278 0.5992 0.6148 0.666
Eco-Driving Inflow 0.4061 0.2229 0.5077 0.5513 0.5299 0.5172 0.5528
Eco-Driving Green Phase 0.385 0.4228 0.4724 0.4697 0.4678 0.4985 0.5027
AA-Ring-Acc Hold Duration 0.8362 0.9209 0.9306 0.9225 0.9136 0.9382 0.9552
AA-Ring-Vel Hold Duration 0.9589 0.972 0.9819 0.9819 0.9819 0.9818 0.9822
AA-Ramp-Acc Hold Duration 0.4276 0.5158 0.6750 0.6821 0.6882 0.6964 0.7111
AA-Ramp-Vel Hold Duration 0.5473 0.5034 0.7170 0.7386 0.6918 0.7273 0.7686
Average 0.6778 0.6017 0.7600 0.7666 0.7563 0.7641 0.7844
@@
We consider three traffic control benchmarks. First, while most traffic lights still operate on fixed schedules, RL can be used to design adaptive (1) Traffic signal control to optimize traffic [8, 24]. However, considering that every intersection is different, challenges persist in generalizing across intersection configurations [19]. Given the significant portion of greenhouse gas emissions in the United States due to transportation [11], the second traffic domain is (2) Dynamic eco-driving at signalized intersections, which concerns learning energy-efficient driving strategies in urban settings. DRL-based eco-driving strategies have been developed [13, 47, 18] but still experience difficulties in generalization. Our final traffic domain is (3) Advisory autonomy, in which real-time speed or acceleration advisories guide human drivers to act as vehicle-based traffic controllers in mixed traffic environments [40, 7, 15]. The context space $X$ is discretized into $N=\{50,50,40\}$ contexts for the three domains, respectively. In Appendix A.4, we provide details about our experiments.
@@
Results. Table 1 and Figure 5 summarize the results. Notably, the Oracle far outperforms the standard baselines (independent and multi-task training), indicating the potential for transfer learning and intelligent training of multiple models, respectively. MBTL rapidly approaches the Oracle within $\approx 10$ transfer steps, indicating that the GP effectively models the training performance and linear generalization gap models the generalization performance. It is important to note that multi-task RL studies commonly consider Independent training as a strong baseline due to its avoidance of negative transfer and other training instability issues. Indeed, independent training often (but not always) outperforms multi-task training in our experiments. Yet, our experiments show that MBTL outperforms both independent and multi-task baselines and matches their performance with up to 30x improved sample efficiency. Among the multi-policy baselines, MBTL often outperforms the heuristic multi-policy baselines, indicating the value of adaptively selecting source tasks based on feedback. The multi-policy baselines, such as random, equidistant, and greedy strategy, also generally outperform Independent and Multi-task, indicating the general value of multi-policy training for solving CMDPs. More results are provided in Appendix A.4.

### v4

@@ 4.1 Modeling assumptions
[…]
@@ 4.2 Bayesian optimization
Gaussian process (GP) regression. Within the framework of BO, we model the source training performance ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\forall x\in X\setminus x_{1:k}}$ using Gaussian process (GP) regression. Specifically, the function ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ is assumed to follow a GP $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)\sim\mathcal{GP}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m(x)},k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})))$, where $m(x)$ is mean and $k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ is the covariance function, representing the expected product of deviations of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{\tilde{x}},}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ from their respective means. Let $D_{k-1}$ denote the data observed up to iteration $k-1$, consisting of the pairs $\{(x_{i},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{i},x_{i})})\}_{i=1,...,k-1}$. The estimated performance $\hat{J}_{k}$ after querying $k-1$ samples is updated as more samples are obtained. The posterior prediction of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}_{k}}$ at a new point $x$, given the data $D_{k-1}$ and previous inputs ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}}$, is normally distributed as $P(\hat{J}_{k}\mid D_{k-1})=\mathcal{N}(\mu_{k}(x),\sigma_{k}^{2}(x))$. $\mu_{k}(x)$ and $\sigma_{k}^{2}(x)$ are defined as $\mu_{k}(x)=\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)]+\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{y}$ and $\sigma_{k}^{2}(x)=k(x,x)-\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{k}$, with $\mathbf{k}$ being the vector of covariances between $x$ and each $x_{i}$ in the observed data, i.e., $\mathbf{k}=[k(x,x_{1}),\ldots,k(x,x_{k-1})]$, and $\mathbf{K}$ is the covariance matrix for the observed inputs, defined as $\mathbf{K}=[k(x_{i},x_{j})]_{1\leq i,j\leq k-1}$. This enables the GP to update its beliefs about the posterior prediction with every new observation, progressively improving the estimation.
@@
$a(x;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}})=\mathbb{E}_{x^{\prime}\in X}[[\mu_{k-1}(x)+\beta_{k}^{1/2}\sigma_{k-1}(x)-\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{{1:k-1}},x^{\prime}})]_{+}]$ (4)
@@ 4.3 Regret analysis
We define the search space $X_{k}$ at iteration $k$ as a subset of $X$, with each element $x^{\prime}\in X_{k}$, such that $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1},x^{\prime}})\leq\hat{J}(\pi_{x_{k}},x^{\prime}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x_{k}},x^{\prime}})$.
@@
For a given $\delta^{\prime}\in(0,1)$ and scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta^{\prime})$, the cumulative regret $R_{K}$ is bounded by $\sqrt{C_{1}\beta_{K}\gamma_{K}\sum_{k=1}^{K}\left(\frac{|X_{k}|}{|X|}\right)^{2}}$ with probability at least $1-\delta^{\prime}$.
@@ 5.1 Setup
Baselines. We consider two types of baselines when evaluating our proposed algorithm: canonical and multi-policy. The canonical baselines are selected to validate multi-policy training; the multi-policy training baselines are heuristic methods designed to validate the Bayesian optimization approach. The canonical baselines include: (1) Independent training, which involves independently training separate models on each task; and (2) Multi-task RL, where a single “universal” context-conditioned policy is trained for all tasks. The multi-policy baselines include: (3) Random selection, where each training task is chosen uniformly at randomly; (4) Greedy strategy, which greedily selects the next source task based on Assumption 1 and fixed training performance; and (5) Sequential Oracle transfer, which has access to generalized performance corresponding to policies for all tasks (including those not yet selected) and uses that information to greedily select the best source task.
Proposed method. In early iterations of BO, GP lacks sufficient observations (often just one or two) and thus relies heavily on its prior. To mitigate this, we incorporate a brief warm-up phase in MBTL. Specifically, we collect three additional data points using a simple greedy approach that selects tasks with the worst observed performance so far (inspired by [7]). After this initialization, the method switches to full BO for source-task selection. We use the scaling factor of $\beta_{k}=2\log(|X|k^{2})$.
DRL algorithms and performance measure. We utilize Deep Q-Networks (DQN) for discrete action spaces [29] and Trust Region Policy Optimization (TRPO) [36] or Proximal Policy Optimization (PPO) [37] for continuous action spaces. For statistical reliability, we run each experiment three times with different random seeds. We evaluate our method by the average performance across all $N$ target tasks after training up to $K=15$ source tasks or the number of source tasks needed to achieve a certain level of performance. We employ min-max normalization of the rewards for each task, and we provide comprehensive details about our model in Appendix A.4.1.
@@ 5.2 Traffic benchmark experiments
Domain Context Variation Independent Multi-task Random Greedy Ours Sequential
Number of Trained Models $N$ $1$ $K$ $K$ $K$ $N$
Traffic Signal Road Length 0.9409 0.8242 0.9366 0.9349 0.9409 0.9432
Traffic Signal Inflow 0.8646 0.8319 0.8699 0.8682 0.8729 0.8773
Traffic Signal Speed Limit 0.8857 0.6083 0.8872 0.8874 0.8866 0.8877
Eco-Driving Penetration Rate 0.5260 0.1945 0.6212 0.5992 0.6519 0.6668
Eco-Driving Inflow 0.4061 0.2229 0.5077 0.5299 0.5356 0.5531
Eco-Driving Green Phase 0.3850 0.4228 0.4724 0.4678 0.4932 0.5058
AA-Ring-Acc Hold Duration 0.8362 0.9219 0.9307 0.9021 0.9329 0.9567
AA-Ring-Vel Hold Duration 0.9589 0.9688 0.9820 0.9819 0.9820 0.9822
AA-Ramp-Acc Hold Duration 0.4276 0.5374 0.6599 0.6570 0.6282 0.7120
AA-Ramp-Vel Hold Duration 0.5473 0.5257 0.7210 0.6461 0.7426 0.7691
Average 0.6778 0.6059 0.7589 0.7474 0.7667 0.7854
@@
We consider three traffic benchmarks. First, while most traffic lights still operate on fixed schedules, RL can be used to design adaptive (1) Traffic signal control to optimize traffic [8, 24]. However, considering that every intersection is different, challenges persist in generalizing across intersection configurations [19]. Given the significant portion of greenhouse gas emissions in the United States due to transportation [11], the second traffic domain is (2) Dynamic eco-driving at signalized intersections, which concerns learning energy-efficient driving strategies in urban settings. DRL-based eco-driving strategies have been developed [13, 48, 18] but still experience difficulties in generalization. Our final traffic domain is (3) Advisory autonomy, in which real-time speed or acceleration advisories guide human drivers to act as vehicle-based traffic controllers in mixed traffic environments [41, 7, 15]. The context space $X$ is discretized into $N=\{50,50,40\}$ contexts for the three domains, respectively. In Appendix A.4, we provide details about our experiments.
@@
Results. Table 1 and Figure 5 summarize the results. Notably, the Oracle far outperforms the standard baselines (independent and multi-task training), indicating the potential for transfer learning and intelligent training of multiple models, respectively. MBTL rapidly approaches the Oracle within $\approx 10$ transfer steps, indicating that the GP effectively models the training performance and linear generalization gap models the generalization performance. It is important to note that multi-task RL studies commonly consider Independent training as a strong baseline due to its avoidance of negative transfer and other training instability issues. Indeed, Independent training often (but not always) outperforms multi-task training in our experiments. Yet, our experiments show that MBTL outperforms both Independent and Multi-task baselines and matches their performance with up to 30x improved sample efficiency. Among the multi-policy baselines, MBTL often outperforms the heuristic multi-policy baselines, indicating the value of adaptively selecting source tasks based on feedback. The multi-policy baselines, such as random and greedy strategy, also generally outperform Independent and Multi-task, indicating the general value of multi-policy training for solving CMDPs. More results are provided in Appendix A.4.